# Global decomposition of networks into multiple cores formed by local hubs

### v2

@@ Abstract
Networks are ubiquitous in various fields, representing systems where nodes and their interconnections constitute their intricate structures. We introduce a network decomposition scheme to reveal multiscale core-periphery structures lurking inside, using the concept of locally defined nodal hub centrality and edge-pruning techniques built upon it. We demonstrate that the hub-centrality-based edge pruning reveals a series of breaking points in network decomposition, which effectively separates a network into its backbone and shell structures. Our local-edge decomposition method iteratively identifies and removes locally least important nodes, and uncovers an onion-like hierarchical structure as a result. Compared with the conventional $k$-core decomposition method, our method based on relative information residing in local structures exhibits a clear advantage in terms of discovering locally crucial substructures. Furthermore, we introduce the core-periphery score to properly separate the core and periphery with our decomposition scheme. By extending the method combined with the network community structure, we successfully detect multiple core-periphery structures by decomposition inside each community. Moreover, the application of our decomposition to supernode networks defined from the communities reveals the intricate relation between the two representative mesoscale structures.
@@ I Introduction
Networks have attracted considerable attention across various fields due to their inevitable omnipresence in nature and society [1, 2, 3]. Consisting of nodes for individual objects of our interest and edges connecting them, a network succinctly represents an interacting system, which is a quintessential topic of statistical mechanics [4]. Compared with traditional topics of statistical mechanics, perhaps a notably distinct feature of the recently developed theory of highly heterogeneous or “complex” networks is the existence of a few dominant elements that can govern the entire system. The identification of such important nodes and their disproportionately significant influence have been studied extensively [1, 2, 3, 4]. One noticeable example in terms of both popularity and significance is the number of the nearest neighboring nodes for each node, which is called the degree. It affects a number of key aspects of networks, e.g., the robustness under failures and attacks [5, 6, 7], the epidemic spreading [8], critical phenomena [9], the controllability [10], etc.
@@
Roughly, there are two streaks of research for identifying important nodes with large degrees within a network: one focuses on individual nodes, e.g., by detecting the ones with large degrees or many connections to the rest of the network, usually dubbed as “hubs” [11], and the other identifies a group of important nodes, e.g., by detecting “core” nodes that are well connected to both each other inside and outside the group  [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. In contrast to the simplest concept of hub nodes by globally counting their neighbors, when it comes to the core node groups, individual nodes’ degree relative to their peers in the group is also crucial. The latter is precisely captured by the measure called “hub centrality” introduced in the series of previous works [23, 7] by some of the authors of this paper, to detect such local hubs in the context of game theory and cascading failure. The hub centrality, defined as the normalized rank of each node within the node group composed of the node itself and its nearest neighbors in terms of degree, successfully identifies locally important nodes. Both global and local hubs play profound roles across various dynamical systems, such as cascading failure [7, 24, 25], disease spreading [12], vaccination [26], and evolutionary game theory [27], according to their unique characteristics.
@@
In this regard, we would like to point out that most of conventional methods to detect core nodes in networks [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] almost exclusively utilizes the concept of global connections in terms of degree, without enough consideration of nodes’ relative position within their neighbor groups. Well-known examples include the decomposition of networks based on degrees, or the k-core decomposition, which iteratively removes nodes with fewer than k connections until only nodes with at least k connections remain [12, 13]. Another related decomposition is the identification of the core-periphery structure [14, 15, 16, 17, 18, 19, 20, 21, 22] by detecting the core nodes with statistically dense connections to the entire network and treating the rest as periphery. For all of these approaches, one simply takes a degree of a node as a face value without consideration of its aforementioned relative position. When there are a variety of local groups in heterogeneous sizes, however, as many real-world networks would actually be so, a particular degree value can make a locally strong hub that governs the entire dynamical property near the hub belonging to a small group, while the same degree value may correspond to a mediocre node belonging to a much larger group. In other words, the mixture of heterogeneous degree distributions [2] and heterogeneous locally dense structures or “communities” [28, 29, 30] requires a meticulous approach to decomposing a network using nodes’ degree as a tool.
@@ II.1 Edge pruning and cusp point
Figure 1: (a) The giant component size $G$ as a function of the fraction $p$ of removed edges. The network is a collaboration network in the field of computational geometry. (b)–(c) The fraction $e_{0}$ of edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ and the fraction $n_{0}$ of nodes with $h=0$ as a function of the cusp point $p_{c}$ and the giant component size $g$ with zero hub-centrality nodes removed, respectively. Each point represents the outcome of each individual network listed in Table 1. The solid lines in (b) and (c) represent $e_{0}=p_{c}$ and $n_{0}=1-g$, respectively.
@@
where $\psi_{i}$ is a certain property of node $i$. In this paper, we try the degree representing the global connectivity and the hub centrality from the normalized local rank in connectivity [23, 7] proposed by some of the authors of this paper. The hub centrality of a node is the fraction of its neighbors with lower degrees than the node [23, 7]. In other words, we first examine the effects of three types of edge importance for pruning edges for network decomposition: the edge betweenness centrality ($C_{B}$[1, 2, 3] by directly setting $\mathscr{E}^{C_{B}}_{ij}\equiv C_{B}(i,j)$, the degree product ($k\mbox{-}\mathrm{P}$) by setting $\mathscr{E}^{k\mbox{-}\mathrm{P}}_{ij}=\psi_{i}\psi_{j}$ where $\psi_{i}\equiv k_{i}$ (the degree of node $i$), and hub-centrality product ($h\mbox{-}\mathrm{P}$) by setting $\mathscr{E}^{h\mbox{-}\mathrm{P}}_{ij}=\psi_{i}\psi_{j}$ where $\psi_{i}\equiv h_{i}$ (the hub centrality of node $i$).
In our edge-pruning strategy, we calculate the set of edge importance $\{\mathscr{E}_{ij}\}$ in the original network and use it throughout the process until the end. We do not recalculate the hub centrality and its product during the edge removal process, as it is not our main interest to investigate the modified structure itself  [^1] Another side effect of such recalculation is that the whole process would depend heavily on the order of removed edges with the same value of $\mathscr{E}_{ij}$, which leads to much more complicated situations in practice that requires ensemble average over multiple realizations.; rather, we would like to extract more central parts in terms of the edge importance in the original structure as we proceed. We implement repeated edge pruning  [^2] We suggest using the heap queue structure in numerical simulations for this., which removes the least important edges at each time step. We first examine a natural measure to characterize the edge-pruning process, which is the relative size $G=|\mathcal{G}|/N$ of giant (i.e., the largest connected) component $\mathcal{G}$ with respect to the original network size $N$, as a function of the fraction $p$ of removed edges. To focus on the fragmentation caused by the edge-pruning process, for all of our numerical studies, we use the giant component of the original networks (the sizes of which are listed inside the parentheses in Table 1) in the beginning.
@@
A typical example is the case of the collaboration network in the field of computational geometry (CB) [37] (see Table 1 for basic statistics) shown in Fig. 1(a). In the case of $k\mbox{-}\mathrm{P}$, $G$ decreases monotonically. For $C_{B}$, the value of $G$ remains high when $p\lesssim 0.4$, but $G$ decreases rapidly once $p\gtrsim 0.4$. Notably, there is a cusp-like point at $p\approx 0.8$, which is the point where the rate of change in $G(p)$ qualitatively changes. For $h\mbox{-}\mathrm{P}$, it shows a more prominent cusp point at $p=p_{c}\approx 0.5$, and the slope of $G(p)$ suddenly drops almost to zero when $p$ crosses the value $p_{c}$. As we will discuss later, in the case of $h\mbox{-}\mathrm{P}$, a significant structural change in the network occurs at the cusp point, whereas we do not observe such a significant change for $C_{B}$. Other networks listed in Table 1 also have the clear-cut cusp point when the edges are removed by $h\mbox{-}\mathrm{P}$ and show qualitatively similar behavior to Fig. 1(a). This seemingly puzzling behavior is, in fact, simply explained, with the hint from the observation of the fraction $e_{0}$ of edges with exactly null importance, i.e., the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$, which are precisely the edges connecting at least one node with the locally smallest degree: $h_{i}=0$ by the product rule. As shown in Fig. 1(b), for all of the four real networks we examine, the $e_{0}$ value coincides with the cusp point $p_{c}$ for each network. This implies that the transition between $p<p_{c}$ and $p>p_{c}$ corresponds to the part where the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ are removed (by our edge-pruning rule, they must be removed first) versus the part where the rest of the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}>0$ are removed.
@@
Moreover, for any node $i$ with $h_{i}=0$, by definition, all of its neighboring nodes should have larger degrees than $k_{i}$, so $\mathscr{E}^{h\mbox{-}\mathrm{P}}_{ij}=0$ for any node $j$ connected to $i$. As a result, all of the edges connected to $i$ will be removed during the $p<p_{c}$ pruning process. Conversely, as already stated, each of the edge-removal processes for $p<p_{c}$ involves such a node with $h_{i}=0$. Therefore, the edge-pruning process for $p<p_{c}$ exactly corresponds to the process of pinpointing the nodes with $h_{i}=0$ and removing all of their edges. This surgical removal effectively isolates those locally least important nodes and leaves the rest of the network as a new giant component precisely at $p=p_{c}$. In principle, it is possible for a node with nonzero hub centrality to be isolated as a result of the removal of all of its neighbors with $h=0$ even for $p<p_{c}$  [^3] One can imagine a following (extreme) case: assume a node with $h>0$ connected to a number of other nodes, all of which have $h=0$ and are connected to very large and dense clique-like subnetworks on the other side than the original node itself. In that case, by the elimination of the neighboring nodes, the original node will be isolated. This effect is reflected in the actual appearance of $g\equiv G(p=p_{c})<1-n_{0}$ in Fig. 2(d), but actually it is quite rare (usually less than $0.1\%$ of nodes in our empirical networks).. However, one can check that the relation $g\equiv G(p=p_{c})=1-n_{0}$ holds (under the assumption of the absence of such cases; in general, $g<1-n_{0}$ caused by the possibility of the aforementioned “casualty” node with $h>0$) for our empirical networks, where $n_{0}$ is the fraction of the nodes with $h_{i}=0$ and $g$ is the fraction of the giant component remaining at $p=p_{c}$, from Fig. 1(c). In other words, the separation of zero-hub-centrality nodes does not cause the noticeable separation of other nodes with nonzero hub centrality from the giant component up to $p=p_{c}$, and most nodes with $h>0$ belong to the new giant component at $p=p_{c}$.
@@
From these observations, we deduce that a network composed of a single giant component harbors a one-step deeper-level giant component formed by positive hub-centrality nodes inside and zero hub-centrality nodes (along with a negligible fraction of positive hub-centrality nodes) attached to it outside. In terms of transport property, if we choose two different zero hub-centrality nodes, as each of them is likely to have a neighbor with $h>0$, there is usually at least a path only through nonzero hub-centrality nodes, which is reminiscent of the backup-pathway-based notion of core-periphery [42, 43] and highlights the role of the giant component at $p=p_{c}$ as a structural and dynamical backbone. On that threshold, the network is divided into a backbone composed of nodes with $h>0$ and a shell mostly composed of nodes with $h=0$.
@@ II.2 Local-edge decomposition and node hierarchy
Figure 2: (a)–(c) The giant component size $G$ as a function of the fraction $p$ of removed edges based on the hub-centrality product ($h\mbox{-}\mathrm{P}$). (a) The original collaboration network used in Fig. 1. (b) The backbone of the original network. (c) The backbone of the original network’s backbone. (d) The relative fraction $g$ of the giant component size with zero hub-centrality nodes removed as a function of the fraction $n_{0}$ of zero hub-centrality nodes in decomposed networks. The solid line represents $n_{0}=1-g$, which implies the set of zero hub-centrality nodes $\approx$ the shell at each decomposition level.
@@
Figure 2(d) demonstrates that our earlier argument holds well; $n_{0}\approx 1-g$ (for $p<p_{c}$) and the edges are eliminated if and only if $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ at each level. As a result, all of the $(n_{0},g)$ pairs at different levels fall on top of the $n_{0}=1-g$ line. The other networks listed in Table 1 also show qualitatively similar results to Fig. 2. The final level is composed of nodes with the same degree and the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ (no node’s degree can be smaller than any other nodes’ degree), so the whole process is naturally terminated by the elimination of the entire edges altogether along with all of the nodes; generally, it forms a clique in real-world networks according to our observation. Therefore, we conclude that a network can be decomposed as an onion-like structure [31], wherein we extract the core of the onion by removing locally least important nodes iteratively through LED.

### v3

@@ Abstract
Networks are ubiquitous in various fields, representing systems where nodes and their interconnections constitute their intricate structures. We introduce a network decomposition scheme to reveal multiscale core-periphery structures lurking inside, using the concept of locally defined nodal hub centrality and edge-pruning techniques built upon it. We demonstrate that the hub-centrality-based edge pruning reveals a series of breaking points in network decomposition, which effectively separates a network into its backbone and shell structures. Our local-edge decomposition method iteratively identifies and removes locally least connected nodes, and uncovers an onion-like hierarchical structure as a result. Compared with the conventional $k$-core decomposition method, our method based on relative information residing in local structures exhibits a clear advantage in terms of discovering locally crucial substructures. As an application of the method, we present a scheme to detect multiple core-periphery structures and the decomposition of coarse-grained supernode networks, by combining the method with the network community detection.
@@ I Introduction
Networks have attracted considerable attention across various fields due to their omnipresence in nature and society [1, 2, 3]. Consisting of nodes for individual objects of our interest and edges connecting them, a network succinctly represents an interacting system, which is a quintessential topic of statistical mechanics [4]. Compared with traditional topics of statistical mechanics, perhaps a notably distinct feature of the recently developed theory of highly heterogeneous or “complex” networks is the existence of a few dominant elements that can govern the entire system. The identification of such important nodes and their disproportionately significant influence have been studied extensively [1, 2, 3, 4]. One noticeable example in terms of both popularity and significance is the number of neighboring nodes for each node, which is called the degree. It affects a number of key aspects of networks, e.g., the robustness under failures and attacks [5, 6, 7], the epidemic spreading [8], critical phenomena [9], the controllability [10], etc.
@@
Roughly, there are two streaks of research for identifying important nodes with large degrees within a network: one focuses on individual nodes, e.g., by detecting the ones with large degrees or many connections to the rest of the network, usually dubbed as “hubs” [11], and the other identifies a group of important nodes, e.g., by detecting “core” nodes that are well connected to both each other inside and outside the group  [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. In contrast to the simplest concept of hub nodes by globally counting their neighbors, when it comes to the core node groups, individual nodes’ degree relative to their peers in the group is also crucial. The latter is precisely captured by the measure called “hub centrality” introduced in the series of previous works [23, 7] by some of the authors of this paper, to detect such local hubs in the context of game theory and cascading failure. The hub centrality, defined as the normalized rank of each node within the node group composed of the node itself and its neighbors in terms of degree, successfully identifies locally important nodes. Both global and local hubs play profound roles across various dynamical systems, such as cascading failure [7, 24, 25], disease spreading [12], vaccination [26], and evolutionary game theory [27], according to their unique characteristics.
@@
In this regard, we point out that most of conventional methods to detect core nodes in networks [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] almost exclusively utilizes the concept of global connections in terms of degree, without enough consideration of nodes’ relative position within their neighbor groups. Well-known examples include the decomposition of networks based on degrees, or the k-core decomposition, which iteratively removes nodes with fewer than k connections until only nodes with at least k connections remain [12, 13]. Another related decomposition is the identification of the core-periphery structure [14, 15, 16, 17, 18, 19, 20, 21, 22] by detecting the core nodes with statistically dense connections to the entire network and treating the rest as periphery. For all of these approaches, one simply takes a degree of a node as a face value without consideration of its aforementioned relative position. When there are a variety of local groups in heterogeneous sizes, however, as many real-world networks would actually be so, a particular degree value can make a locally strong hub that governs the entire dynamical property near the hub belonging to a small group, while the same degree value may correspond to a mediocre node belonging to a much larger group. In other words, the mixture of heterogeneous degree distributions [2] and heterogeneous locally dense structures or “communities” [28, 29, 30] requires a meticulous approach to decomposing a network using nodes’ degree as a tool.
@@ II.1 Edge pruning and cusp point
Figure 1: (a) Relative size $G$ of the giant component as a function of the fraction $p$ of removed edges. The network is a collaboration network in the field of computational geometry. (b)–(c) Fraction $e_{0}$ of edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ and the fraction $n_{0}$ of nodes with $h=0$ as a function of the cusp point $p_{c}$ and the giant component size $g$ with zero hub-centrality nodes removed, respectively. Each point represents the outcome of each individual network listed in Table 1. The solid lines in (b) and (c) represent $e_{0}=p_{c}$ and $n_{0}=1-g$, respectively.
@@
where $\psi_{i}$ is a certain property of node $i$. In this paper, we try the degree representing the global connectivity and the hub centrality from the normalized local rank in connectivity [23, 7] proposed by some of the authors of this paper. The hub centrality of a node is the fraction of its neighbors with strictly lower degrees than the node itself [23, 7], given by
$h_{i}=\frac{N_{<k_{i}}}{k_{i}}\,,$ (2)
where $N_{<k_{i}}$ is the number of neighbors of node $i$ with lower degrees than the node $i$ itself, and the denominator $k_{i}$ is the degree of the node $i$ for normalization $0\leq h_{i}\leq 1$. We first try examining the effects of three types of edge importance for pruning edges in network decomposition, starting from widely used centralities: degree and betweenness centrality, in comparison to the hub centrality. We use the edge betweenness centrality ($C_{B}$[1, 2, 3] by directly setting $\mathscr{E}^{C_{B}}_{ij}\equiv C_{B}(i,j)$, the degree product ($k\mbox{-}\mathrm{P}$) by setting $\mathscr{E}^{k\mbox{-}\mathrm{P}}_{ij}=\psi_{i}\psi_{j}$ where $\psi_{i}\equiv k_{i}$, and hub-centrality product ($h\mbox{-}\mathrm{P}$) by setting $\mathscr{E}^{h\mbox{-}\mathrm{P}}_{ij}=\psi_{i}\psi_{j}$ where $\psi_{i}\equiv h_{i}$.
In our edge-pruning strategy, we calculate the set of edge importance $\{\mathscr{E}_{ij}\}$ in the original network and use it throughout the process until the end. We do not recalculate the hub centrality and its product during the edge removal process, as it is not our main interest to investigate the modified structure itself  [^1] Another side effect of such recalculation is that the whole process would depend heavily on the order of removed edges with the same value of $\mathscr{E}_{ij}$, which leads to much more complicated situations in practice that requires ensemble average over multiple realizations.; rather, we extract more central parts in terms of the edge importance in the original structure as we proceed. We implement repeated edge pruning  [^2] We suggest using the heap queue structure in numerical simulations for this., which removes the least important edges at each time step. We first examine a natural measure to characterize the edge-pruning process, which is the relative size $G=|\mathcal{G}|/N$ of giant (i.e., the largest connected) component $\mathcal{G}$ with respect to the original network size $N$, as a function of the fraction $p$ of removed edges. To focus on the fragmentation caused by the edge-pruning process, for all of our numerical studies, we use the giant component of the original networks (the sizes of which are listed inside the parentheses in Table 1) in the beginning.
@@
A typical example is the case of the collaboration network in the field of computational geometry (CB) [37] (see Table 1 for basic statistics) shown in Fig. 1(a); qualitatively similar results are observed across other networks listed in Table 1, along with those from other networks not listed there. In the case of $k\mbox{-}\mathrm{P}$, $G$ decreases monotonically. For $C_{B}$, the value of $G$ remains high when $p\lesssim 0.4$, but $G$ decreases rapidly once $p\gtrsim 0.4$. Notably, there is a cusp-like point at $p\approx 0.8$, which is the point where the rate of change in $G(p)$ qualitatively changes. For $h\mbox{-}\mathrm{P}$, it shows a more prominent cusp point at $p=p_{c}\approx 0.5$, and the slope of $G(p)$ suddenly drops almost to zero when $p$ crosses the value $p_{c}$. As we will discuss later, in the case of $h\mbox{-}\mathrm{P}$, a significant structural change in the network occurs at the cusp point, whereas we do not observe such a significant change for $C_{B}$. Other networks listed in Table 1 also have the clear-cut cusp point when the edges are removed by $h\mbox{-}\mathrm{P}$ and show qualitatively similar behavior to Fig. 1(a). This seemingly puzzling behavior is, in fact, simply explained, with the hint from the observation of the fraction $e_{0}$ of edges with exactly null importance, i.e., the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$, which are precisely the edges connecting at least one node with the locally smallest degree: $h_{i}=0$ by the product rule. As shown in Fig. 1(b), for all of the four real networks we examine, the $e_{0}$ value coincides with the cusp point $p_{c}$ for each network. This implies that the transition between $p<p_{c}$ and $p>p_{c}$ corresponds to the part where the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ are removed (by our edge-pruning rule, they must be removed first) versus the part where the rest of the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}>0$ are removed. Note that we have tried other types of centrality measures including closeness centrality, but none of them shows this peculiar behavior.
@@
Moreover, for any node $i$ with $h_{i}=0$, by definition, all of its neighboring nodes should have larger degrees than $k_{i}$, so $\mathscr{E}^{h\mbox{-}\mathrm{P}}_{ij}=0$ for any node $j$ connected to $i$. As a result, all of the edges connected to $i$ will be removed during the $p<p_{c}$ pruning process. Conversely, as already stated, each of the edge-removal processes for $p<p_{c}$ involves such a node with $h_{i}=0$. Therefore, the edge-pruning process for $p<p_{c}$ exactly corresponds to the process of pinpointing the nodes with $h_{i}=0$ and removing all of their edges. This surgical removal effectively isolates those locally least connected nodes and leaves the rest of the network as a new giant component precisely at $p=p_{c}$. In principle, it is possible for a node with nonzero hub centrality to be isolated as a result of the removal of all of its neighbors with $h=0$ even for $p<p_{c}$  [^3] One can imagine a following (extreme) case: assume a node with $h>0$ connected to a number of other nodes, all of which have $h=0$ and are connected to very large and dense clique-like subnetworks on the other side than the original node itself. In that case, by the elimination of the neighboring nodes, the original node will be isolated. This effect is reflected in the actual appearance of $g\equiv G(p=p_{c})<1-n_{0}$ in Fig. 2(d), but actually it is quite rare (usually less than $0.1\%$ of nodes in our empirical networks).. However, one can check that the relation $g\equiv G(p=p_{c})=1-n_{0}$ holds (under the assumption of the absence of such cases; in general, $g<1-n_{0}$ caused by the possibility of the aforementioned “casualty” node with $h>0$) for our empirical networks, where $n_{0}$ is the fraction of the nodes with $h_{i}=0$ and $g$ is the fraction of the giant component remaining at $p=p_{c}$, from Fig. 1(c). In other words, the separation of zero-hub-centrality nodes does not cause the noticeable separation of other nodes with nonzero hub centrality from the giant component up to $p=p_{c}$, and most nodes with $h>0$ belong to the new giant component at $p=p_{c}$.
@@
From these observations, we deduce that, except for peculiar substructures such as linear chains attached to the component  [^4] Even in the case of a long linear chain, one can easily show that only a single node with $h>0$ right next to the dead end is removed per chain and all of the other nodes in the chain has the value $h=0$ anyway., a network composed of a single giant component harbors a one-step deeper-level giant component formed by positive hub-centrality nodes inside and zero hub-centrality nodes attached to it outside. Simply speaking, inside nodes maintain the connectivity even when outside nodes are removed. In terms of transport property, if we choose two different zero hub-centrality nodes (outside nodes), as each of them has connections to nonzero hub-centrality nodes (inside nodes), there is at least a path only through nodes with $h>0$, which is reminiscent of the backup-pathway-based notion of core-periphery [43, 44] and highlights the role of the giant component at $p=p_{c}$ as a structural and dynamical backbone. On that threshold, the network is divided into a backbone composed of nodes with $h>0$ and a shell mostly composed of nodes with $h=0$.
@@ II.2 Local-edge decomposition and node hierarchy
Figure 2: (a)–(c) Relative size $G$ of the giant component as a function of the fraction $p$ of removed edges based on the hub-centrality product ($h\mbox{-}\mathrm{P}$). (a) The original collaboration network used in Fig. 1. (b) Backbone of the original network. (c) Backbone of the original network’s backbone. (d) Relative fraction $g$ of the giant component size with zero hub-centrality nodes removed as a function of the fraction $n_{0}$ of zero hub-centrality nodes in decomposed networks. The solid line represents $n_{0}=1-g$, which implies the set of zero hub-centrality nodes $\approx$ the shell at each decomposition level.
@@
Figure 2(d) demonstrates that our earlier argument holds well; $n_{0}\approx 1-g$ (for $p<p_{c}$) and the edges are eliminated if and only if $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ at each level. As a result, all of the $(n_{0},g)$ pairs at different levels fall on top of the $n_{0}=1-g$ line. The other networks listed in Table 1 also show qualitatively similar results to Fig. 2. The final level is composed of nodes with the same degree and the edges with $\mathscr{E}^{h\mbox{-}\mathrm{P}}=0$ (no node’s degree can be smaller than any other nodes’ degree), so the whole process is naturally terminated by the elimination of the entire edges altogether along with all of the nodes; generally, it forms a clique in real-world networks according to our observation. Therefore, we conclude that a network can be decomposed as an onion-like structure [31], wherein we extract the core of the onion by removing locally least connected nodes iteratively through LED.